# Vectorized Computation of Euler Characteristic Functions and Transforms

### v2

@@ 3 Computation of WECFs
We fix a weighted simplicial complex $\mathbb{K}=(K,\omega)$ and a set of vertex filters $\{f_{0},f_{1},\ldots,f_{m-1}\}$ for the entirety of this section. To simplify notation, we extend each $f_{p}:K_{0}\to\mathbb{R}$ to $K$ by defining $f_{p}^{\max}:K\to\mathbb{R}$ as
@@
We fix enumerations $K_{0}=\{v_{0},v_{1},\ldots,v_{k_{0}}\}$ and $K_{i}=\{\sigma_{0}^{i},\sigma_{1}^{i},\ldots,\sigma_{k_{i}}^{i}\}$ for each $i\geq 1$.
@@
Algorithm 1 ultimately produces a tensor WECFs with shape $(m,\texttt{NumVals})$ and values
$\texttt{WECFs}[p,q]=\operatorname{WECF}_{f_{p}}\big(\beta(q)\big).$
@@
The operation $\texttt{ScatterAdd}(\texttt{VIndices}^{T},\texttt{VertexWeights})$ is then applied to diffWECFs, resulting in adding to diffWECFs, the tensor $D_{0}$ with shape $(m,\texttt{NumVals})$ and values
@@
Next, inside the for-loop, fix an $i\in\{1,2,\ldots\dim(K)\}$. The tensor SimplexVertices has rows indexed by $K_{i}$ with the entries in each row corresponding to the vertex sets of simplices in $K_{i}$. Therefore, the tensor MaxSimpIndices has shape $(k_{i},m)$ and values
@@
The operation $\texttt{ScatterAdd}(\texttt{MaxSimpIndices}^{T},(-1)^{i}*\texttt{SimplexWeights})$ updates diffWECFs by adding to it, the tensor $D_{i}$ with shape $(m,\texttt{NumVals})$ and values
@@
$\texttt{CumulativeSum}(D_{i})[p,q]=\sum_{\begin{subarray}{c}\sigma\in K_{i}\\ f_{p}^{\max}(\sigma)\leq\beta(q)\end{subarray}}\omega(\sigma).$
@@
$\texttt{WECFs}[p,q]$ $=\sum_{\begin{subarray}{c}v\in K_{0}\\ f_{p}(v)\leq\beta(q)\end{subarray}}\omega(v)+\sum_{i=1}^{\dim(K)}\sum_{\begin{subarray}{c}\sigma\in K_{i}\\ f_{p}^{\max}(\sigma)\leq\beta(q)\end{subarray}}\omega(\sigma)$
$=\sum_{\begin{subarray}{c}\sigma\in K\\ f_{p}^{\max}(\sigma)\leq\beta(q)\end{subarray}}\omega(\sigma)$
@@
The algorithm achieves a theoretically optimal final time and memory complexity of $\Theta(k)$, where $k$ is the number of simplices in $K$. We refer the reader to Appendix B for the full complexity analysis of Algorithm 1.
@@ 4 Experiments
In order to demonstrate the efficacy of our implementation, we evaluate and compare the runtime of Algorithm 1 with existing methods; see Section 2.3 on a variety of datasets. [^7] The implementation of pyECT can be accessed here: https://github.com/compTAG/pyECT. The source code for reproducing experiments (with instructions) is available here: https://github.com/compTAG/pyECT-experiments.
@@ 4.2 Results
We now present the experimental results, measured by the average speedup factor of our method over the existing methods. The average speedup factor is defined as the runtime of the baseline method (e.g., EuCalc, DECT, or FastTopology) divided by the runtime of the proposed method, pyECT, averaged across all images in the dataset.
@@
Figure 2: A comparison for computing on the Fashion-MNIST dataset for WECT computation. Figures (2(a)) and (2(b)) compare pyECT with EuCalc excluding the vectorization time, (2(c)) and (2(d)) compare pyECT with EuCalc including the vectorization time, and (2(e)) and (2(f)) compare pyECT with DECT. CPU implementations of pyECT are shown on the left and GPU implementations are shown on the right. Observe that vertical axes are not consistent across subfigures. In nearly all cases, pyECT outperforms the baseline methods. This behavior is especially pronounced when using the GPU architecture with large numbers of directions and filtration heights.
@@
Figure 3: A comparison for computing on the ImageNet dataset for WECT computation. Figures (3(a)) and (3(b)) compare pyECT with EuCalc excluding the vectorization time, (3(c)) and (3(d)) compare pyECT with EuCalc including the vectorization time, and (3(e)) and (3(f)) compare pyECT with DECT. Note that vertical axes are not consistent across subfigures. In nearly all cases, pyECT outperforms the baseline methods. This behavior is especially pronounced when using the GPU architecture with large numbers of directions and filtration heights. Note that DECT was unable to compute the WECT for large numbers of directions and heights due to memory constraints, resulting in missing values in (3(e)) and (3(f)). However, pyECT was able to compute the WECT in these scenarios.
Speedup results for the comparison of pyECT against DECT on the three-dimensional mesh datasets are presented in Figure 4. Note that experiments were run 10 times on each mesh dataset to generate the average and standard deviation values shown in the plots. For large choices of directions and filtration heights in both datasets, DECT was unable to compute the ECT due to memory constraints. While this results in missing values in all subplots, from the cases where DECT was able to compute the ECT, pyECT outperforms DECT (especially for larger choices of parameters). The memory issues are more pronounced in the Armadillo dataset (due to its larger size). Note that while pyECT demonstrates significant speedups in the GPU implementations, its advantage is even more pronounced in the CPU implementations, highlighting the versatility of pyECT across different hardware architectures.
@@
Figure 5: A comparison of average speedup of pyECT over FastTopology when computing the ECF of multiple datasets. The comparison evaluates both uncompiled and compiled versions of pyECT. Note that vertical axes are not consistent across subfigures. In nearly all cases, pyECT outperforms FastTopology. This behavior is especially pronounced when using the GPU architecture on the three-dimensional 3DCCs dataset.
@@ 5 Discussion
In this work, we present a vectorized framework for computing WECFs (including ECFs and the WECT). This framework allows for efficient computation of these functions using tensor operations, which are well-suited for modern GPU hardware. Furthermore, our method works in full generality for weighted simplicial complexes (or cubical complexes) of arbitrary dimension with no structural constraints. Our framework is implemented in a publicly available Python package called pyECT. Experimentally, pyECT demonstrates a significant speedup over existing methods across a variety of image datasets. The implementation is also incredibly concise as the core vectorized operations leverage PyTorch, a well-maintained and optimized library for tensor computation. The core WECT algorithm in the pyECT implementation is only about $25$ lines of code, which we acknowledge is not a perfect metric, but is a testament to the simplicity of the approach.
The pyECT package also contains a variety of tools for constructing and working with simplicial/cubical complexes. Our package includes fully vectorized methods for building weighted Freudenthal complexes and weighted cubical complexes from images. Additionally, pyECT has dedicated methods for the specific cases of computing ECFs of images and WECTs of weighted geometric simplicial complexes in $\mathbb{R}^{n}$. These two special cases are implemented as PyTorch neural network modules.
Another benefit of the approach presented in this work is that vectorized operations support automatic gradient computations. Potential future work includes exploring the use of this framework in differentiable forms of topological transforms (as in [17]) and, in particular, incorporating these transforms into neural network architectures. Additionally, we believe that this work will encourage further investigation into vectorized methods for computing other topological functions and transforms.
@@ Acknowledgements
We thank Jack Ruder and Jacob Sriraman for their valuable assistance in developing and refining portions of the methodology and code used in this work. Computational efforts were performed on the Tempest High Performance Computing System, operated and supported by University Information Technology Research Cyberinfrastructure (RRID:SCR_026229) at Montana State University.
@@ Appendix A Examples of Tensor Operations
Appendix B Complexity Analysis
The function ComputeWECFs has inputs Complex and NumVals requiring $\Theta(km)$ memory usage. The algorithm first initializes a zero tensor diffWECFs of shape $(m,\texttt{NumVals})$, requiring $\Theta(m\texttt{NumVals})$ memory. The function $\alpha$ (see Equation (6)) is then applied elementwise to Filters, taking $\Theta(km)$ time.
Now, for a fixed dimension $i\in\{1,2,\ldots n\}$, the tensor
$\texttt{SimpIndices}=\texttt{VIndices}[\texttt{SimplexVertices}]$
which takes $\Theta(km)$ time and memory.
Thus, we have a total time and memory complexity of $\Theta(nkm)=\Theta(km)$ after conclusion of the for-loop. Finally, a cumulative sum is applied to diffWECFs, taking $\Theta(m\texttt{NumVals})$ time and using no additional memory. This results in $\Theta(m(k+\texttt{NumVals}))$ for Algorithm 1 as the final time and memory complexity. In practice, $m$ and NumVals are typically held constant. Indeed, this is necessary to achieve consistent vectorizations of data. Holding $m$ and NumVals constant results in a theoretically optimal final time and memory complexity of $\Theta(k)$ for Algorithm 1.

### v3

@@ 3 Computation of WECFs
We fix a weighted simplicial complex $\mathbb{K}=(K,\omega)$ and a set of vertex filters $\{f_{0},f_{1},\ldots,f_{m-1}\}$ for the entirety of this section. To simplify notation, we extend each $f_{p}\colon K_{0}\to\mathbb{R}$ to $K$ by defining $f_{p}^{\max}\colon K\to\mathbb{R}$ as
@@
We fix enumerations $K_{0}=\{v_{0},v_{1},\ldots,v_{k_{0}-1}\}$ and $K_{i}=\{\sigma_{0}^{i},\sigma_{1}^{i},\ldots,\sigma_{k_{i}-1}^{i}\}$ for each $i\geq 1$.
@@
Here, we prove that Algorithm 1 produces the tensor WECFs with shape $(m,\texttt{NumVals})$ and values
$\texttt{WECFs}[p,q]=\operatorname{WECF}_{f_{p}}\big(\beta(q)\big).$ (7)
@@
The operation $\texttt{ScatterAdd}(\texttt{VIndices}^{T},\texttt{VertexWeights})$ is then applied to diffWECFs, resulting in adding to diffWECFs the tensor $D_{0}$ with shape $(m,\texttt{NumVals})$ and values
@@
Next, inside the for-loop, fix an $i\in\{1,2,\ldots,\dim(K)\}$. The tensor SimplexVertices has rows indexed by $K_{i}$ with the entries in each row corresponding to the vertex sets of simplices in $K_{i}$. Therefore, the tensor MaxSimpIndices has shape $(k_{i},m)$ and values
@@
The operation $\texttt{ScatterAdd}(\texttt{MaxSimpIndices}^{T},(-1)^{i}*\texttt{SimplexWeights})$ updates diffWECFs by adding to it the tensor $D_{i}$ with shape $(m,\texttt{NumVals})$ and values
@@
$\texttt{CumulativeSum}(D_{i})[p,q]=\sum_{\begin{subarray}{c}\sigma\in K_{i}\\ f_{p}^{\max}(\sigma)\leq\beta(q)\end{subarray}}(-1)^{i}\omega(\sigma).$
@@
$\texttt{WECFs}[p,q]$ $=\sum_{\begin{subarray}{c}v\in K_{0}\\ f_{p}(v)\leq\beta(q)\end{subarray}}\omega(v)+\sum_{i=1}^{\dim(K)}\sum_{\begin{subarray}{c}\sigma\in K_{i}\\ f_{p}^{\max}(\sigma)\leq\beta(q)\end{subarray}}(-1)^{i}\omega(\sigma)$
$=\sum_{\begin{subarray}{c}\sigma\in K\\ f_{p}^{\max}(\sigma)\leq\beta(q)\end{subarray}}(-1)^{\dim(\sigma)}\omega(\sigma)$
@@
3.1 Complexity Analysis
The input Complex requires $\Theta(km)$ memory usage. The algorithm first initializes a zero tensor diffWECFs of shape $(m,\texttt{NumVals})$, requiring $\Theta(m\texttt{NumVals})$ memory. The function $\alpha$ (see Equation (5)) is then applied elementwise to Filters, taking $\Theta(km)$ time.
Now, for a fixed dimension $i\in\{1,2,\ldots n\}$, the tensor
$\texttt{SimpIndices}=\texttt{VIndices}[\texttt{SimplexVertices}]$
which takes $\Theta(km)$ time and memory.
Thus, we have a total time and memory complexity of $\Theta(nkm)=\Theta(km)$ after conclusion of the for-loop. Finally, a cumulative sum is applied to diffWECFs, taking $\Theta(m\texttt{NumVals})$ time and using no additional memory. This results in $\Theta(m(k+\texttt{NumVals}))$ for Algorithm 1 as the final time and memory complexity. In practice, $m$ and NumVals are typically held constant. Indeed, this is necessary to achieve consistent vectorizations of data. Holding $m$ and NumVals constant results in a final time and memory complexity of $\Theta(k)$ for Algorithm 1.
@@ 4 Experiments
In order to demonstrate the efficacy of our implementation, we evaluate and compare the runtime of Algorithm 1 with existing methods (see Section 2.2) on a variety of datasets. [^7] The implementation of pyECT can be accessed here: https://anonymous.4open.science/r/pyECT-anonymous-902F/. The source code for reproducing experiments (with instructions) is available here: https://anonymous.4open.science/r/pyECT-experiments-anonymous-0E4F/.
@@ 4.2 Results
We now present the experimental results, measured by the average speedup factor of our method over existing methods. The average speedup factor is defined as the runtime of the baseline method (e.g., EuCalc, DECT, or FastTopology) divided by the runtime of the proposed method, pyECT, averaged across all images in the dataset. Across all figures, error bars represent $95\%$ confidence intervals.
@@
Figure 2: A comparison for computing on the Fashion-MNIST dataset for WECT computation. Figures (2(a)) and (2(b)) compare pyECT with EuCalc excluding the vectorization time, (2(c)) and (2(d)) compare pyECT with EuCalc including the vectorization time, and (2(e)) and (2(f)) compare pyECT with DECT. CPU implementations of pyECT are shown on the left and GPU implementations are shown on the right. Vertical axes are not consistent across subfigures. In nearly all cases, pyECT outperforms the baseline methods. This behavior is especially pronounced when using the GPU architecture with large numbers of directions and filtration heights.
@@
Figure 3: A comparison for computing on the ImageNet dataset for WECT computation. Figures (3(a)) and (3(b)) compare pyECT with EuCalc excluding the vectorization time, (3(c)) and (3(d)) compare pyECT with EuCalc including the vectorization time, and (3(e)) and (3(f)) compare pyECT with DECT. Note that vertical axes are not consistent across subfigures. DECT was unable to compute the WECT for large numbers of directions and heights due to memory constraints, resulting in missing values in (3(e)) and (3(f)). However, pyECT was able to compute the WECT in these scenarios.
Speedup results for the comparison of pyECT against DECT on the three-dimensional mesh datasets are presented in Figure 4. Note that experiments were run 10 times on each mesh dataset to generate the average and confidence interval values shown in the plots. For large choices of directions and filtration heights in both datasets, DECT was unable to compute the ECT due to memory constraints. While this results in missing values in all subplots, from the cases where DECT was able to compute the ECT, pyECT outperforms DECT (especially for larger choices of parameters). The memory issues are more pronounced in the Armadillo dataset (due to its larger size). Note that while pyECT demonstrates significant speedups in the GPU implementations, its advantage is even more pronounced in the CPU implementations, highlighting the versatility of pyECT across different hardware architectures.
@@
Next, to verify the efficiency of the computation on the GPU, we compute the ECF of 2D and 3D images with intensity drawn uniformly and measure throughput as in [26, 25]. For 2D images (sizes $2048^{2}$, $4096^{2}$, and $8192^{2}$) with data pre-loaded into GPU memory, uncompiled pyECT achieves 1.111.111.131.13 GPix/s at 1,0001{,}000 thresholds and 2.402.402.522.52 GPix/s at 10,00010{,}000 thresholds, while the compiled version achieves 1.281.281.321.32 GPix/s and 2.522.524.114.11 GPix/s, respectively. For 3D images (sizes $128^{3}$, $256^{3}$, and $512^{3}$), uncompiled pyECT achieves 0.480.480.490.49 GVox/s at 1,0001{,}000 thresholds and 1.051.051.091.09 GVox/s at 10,00010{,}000 thresholds, while the compiled version achieves 0.430.430.560.56 GVox/s and 1.091.091.761.76 GVox/s, respectively.
@@
Figure 5: A comparison of average speedup of pyECT over FastTopology when computing the ECF of multiple datasets. The comparison evaluates both uncompiled and compiled versions of pyECT. Note that vertical axes are not consistent across subfigures.
@@ 5 Discussion
In this work, we present a tensor-based framework for computing WECFs (including ECFs and the WECT). This framework allows for efficient computation of these functions using tensor operations, which are well-suited for modern GPU hardware. Furthermore, our method works in full generality for weighted simplicial complexes (or cubical complexes) of arbitrary dimension. Our framework is implemented in a publicly available Python package called pyECT. Experimentally, pyECT demonstrates a significant speedup over existing methods across a variety of datasets.
The pyECT package also contains a variety of tools for constructing and working with simplicial/cubical complexes. Our package includes fully vectorized methods for building weighted Freudenthal complexes and weighted cubical complexes from images. Additionally, pyECT has dedicated methods for computing ECFs of images and WECTs of weighted geometric simplicial complexes in $\mathbb{R}^{n}$. These two special cases are implemented as PyTorch neural network modules. Future work includes using this framework in differentiable forms of topological transforms (as in [17]) and incorporating these transforms into neural network architectures. Additionally, we believe that this work will encourage further investigation into vectorized methods for computing other topological functions and transforms.
@@ Acknowledgements
We thank Jack Ruder and Jacob Sriraman for their valuable assistance in developing and refining portions of the methodology and code used in this work. Computational efforts were performed on the Tempest High Performance Computing System, operated and supported by the University Information Technology Research Cyberinfrastructure (RRID:SCR_026229) at Montana State University.